Using Stoke’s Theorem evaluate the line integral ∫ ↓c F. dl where C is the ellipse


x²/4 +y²/16 = 1 in the xy - plane and F vector = 2x²i^ +4xj^ +2z²k^.

1
Expert's answer
2022-02-28T13:52:45-0500

"F=2x^2\hat{i}+4x\hat{j}+2z^2\hat{k}"

"dr=dx\hat{i}+dy{j}"

Now stock theorem


"\smallint F.dr=\smallint (2x^2\hat{i}+4x\hat{j}+2z^2\hat{k}).(dx\hat{i}+dy{j})=\frac{2x^3}{3}+4xy"

"\smallint F.dr= \frac{2x^3}{3}+4xy"

At point

P(2,4)

"\smallint F.dr=\frac{2\times2^3}{3}+4\times2\times4"


"\smallint F.dr=\frac{16}{3}+32=\frac{112}{3}=37.33J"